# Blessing of dimension in Bayesian inference on covariance matrices

### v1

@@ Abstract
Bayesian factor analysis is routinely used for dimensionality reduction in modeling of high-dimensional covariance matrices. Factor analytic decompositions express the covariance as a sum of a low rank and diagonal matrix. In practice, Gibbs sampling algorithms are typically used for posterior computation, alternating between updating the latent factors, loadings, and residual variances. In this article, we exploit a blessing of dimensionality to develop a provably accurate pseudo-posterior for the covariance matrix that bypasses the need for Gibbs or other variants of Markov chain Monte Carlo sampling. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach relies on a first-stage singular value decomposition (SVD) to estimate the latent factors, and then defines a jointly conjugate prior for the loadings and residual variances. The accuracy of the resulting pseudo-posterior for the covariance improves with increasing dimensionality. We show that FABLE has excellent performance in high-dimensional covariance matrix estimation, including producing well calibrated credible intervals, both theoretically and through simulation experiments. We also demonstrate the strength of our approach in terms of accurate inference and computational efficiency by applying it to a gene expression data set.
@@ 1 Introduction
Inference on covariance matrices in high-dimensional data is a key focus in many application areas, motivating a rich literature on associated statistical methods. One thread of this literature avoids modeling of the data and instead focuses on high-dimensional covariance matrix estimators under various assumptions on the inherent low-dimensional structure in the data, including (but not limited to) banded covariance (Bickel and Levina,, 2008), low rank structure (Shikhaliev et al.,, 2019), low rank with sparsity (Richard et al.,, 2012), sparse covariance (Bien and Tibshirani,, 2011), and sparse inverse precision matrix estimation (Zhang and Zou,, 2014). Our interest is instead in model-based Bayesian approaches, which have advantages in terms of naturally accommodating complexities in the data and quantifying uncertainty, while having disadvantages in terms of computational efficiency, particularly in implementing Markov chain Monte Carlo (MCMC) algorithms to sample from the posterior distribution.
@@
Our particular focus is on factor models, which express the covariance matrix in a low rank plus diagonal form, and represent a popular probabilistic model-based alternative to principal components analysis (PCA). Factor analysis introduces a lower-dimensional latent factor $\eta_{i}\in\mathbb{R}^{k}$ corresponding to each observation $y_{i}\in\mathbb{R}^{p}$ with $k\ll p$, such that $y_{i}=\Lambda\eta_{i}+\epsilon_{i}$ for a factor loadings matrix $\Lambda\in\mathbb{R}^{p\times k}$ and independent idiosyncratic errors $\epsilon_{i}\sim N_{p}(0,\Sigma)$. Following typical practice, we let the latent factors have independent Gaussian priors $\eta_{i}\sim N_{k}(0,\mathbb{I}_{k})$. Marginalizing out the latent factors leads to $y_{i}\sim N_{p}(0,\Psi)$, with the covariance matrix $\Psi=\Lambda\Lambda^{\top}+\Sigma$. This provides a convenient decomposition of the covariance, which dramatically reduces the number of free parameters required to model the $p\times p$ covariance matrix from $p(p+1)/2$ to $(p-(k-1)/2)k+p$.
@@
There is a vibrant recent literature improving upon and expanding the scope of Bayesian factor analysis (Schiavon et al.,, 2022; De Vito et al.,, 2021; Frühwirth-Schnatter,, 2023; Roy et al.,, 2021; Ma and Liu,, 2022; Bolfarine et al.,, 2022; Xie et al.,, 2022). Even with increasingly rich classes of priors and data types, the canonical approach for posterior computation remains Gibbs samplers that iterate between updating latent factors, factor loadings, residual variances, hyperparameters controlling the hierarchical prior, and other model parameters. This approach is simple to implement in broad model classes, but commonly faces problems with slow mixing, particularly as data dimensionality and complexity increase. Issues with slow mixing and bottlenecks with large datasets have motivated alternative posterior sampling algorithms that apply Hamiltonian Monte Carlo (HMC) after marginalizing out the latent factors, exploiting sufficient statistics and parallel computation to facilitate implementation for large sample sizes (Chandra et al.,, 2023; Man and Culpepper,, 2022). Nonetheless, all current MCMC sampling algorithms for large covariance matrices face substantial computational hurdles as the number of dimensions increases.
@@
The lack of sufficiently scalable MCMC algorithms has motivated a rich literature on developing more computationally efficient algorithms for Bayesian inference in factor models. Variational Bayes approaches (Hansen et al.,, 2023; Wang and Stephens,, 2021) attempt to approximate the posterior but often the accuracy of the approximation is poor, with a tendency to dramatically underestimate posterior uncertainty. A concise review comparing variational approximations and posterior sampling approaches for factor analysis is provided in Foo and Shim, (2021). Maximum a posteriori (MAP) estimation under sparsity priors (Srivastava et al.,, 2017) and expectation-maximization (EM) algorithms (Ročková and George,, 2016; Avalos-Pacheco et al.,, 2022; Zhao et al.,, 2016) provide computationally efficient point estimates of the covariance matrix. Sabnis et al., (2016) develop a distributed computing framework aggregating local estimates of the covariance across different cores. Srivastava et al., (2017); Ročková and George, (2016), and Sabnis et al., (2016) provide theoretical results regarding estimation accuracy and posterior concentration, without providing any guarantees on uncertainty quantification. A key open problem in the literature remains how to obtain an efficient posterior approximation for a high-dimensional covariance that has guarantees in terms of valid frequentist coverage of credible intervals.
@@
Fast algorithms for Bayesian factor analysis capable of scaling efficiently to high-dimensional data typically sacrifice the ability to provide an accurate characterization of uncertainty. The focus of this article is on proposing a simple approach for overcoming this limitation, providing computationally efficient inference on a high-dimensional covariance matrix. With the increasing dimension, we have more and more variables loading on the same small number of latent factors. As a consequence, we obtain a blessing of dimensionality phenomenon allowing us to first pre-estimate the latent factors and then leverage on this point estimate to obtain a pseudo-posterior distribution of the factor loadings $\Lambda$ and error variances $\Sigma$ under conjugate priors on these parameters. This in turn induces a pseudo-posterior distribution for the covariance matrix $\Psi=\Lambda\Lambda^{\top}+\Sigma$. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach completely bypasses MCMC. To ensure that the pseudo-posterior credible intervals for the entries of the covariance are asymptotically normal with guarantees on valid frequentist coverage, we develop a data-adaptive coverage correction (CC); we refer to the adjusted FABLE approach as CC-FABLE.
@@
Fan et al., (2023) also pre-estimate latent factors, but in the context of regression with high-dimensional correlated covariates. In their approach, a first-stage principal components analysis is carried out for the high-dimensional covariates. The results are then used to obtain a modified covariate vector consisting of principal component scores and de-correlated versions of the original covariates. Using this new covariate vector in the second stage high-dimensional regression avoids multicollinearity. This focus, and the corresponding methodology, are substantially different from ours. A blessing of dimensionality phenomenon is also explored in Li et al., (2018), where estimation of the covariance between a subset of the variables of interest is improved by incorporating auxiliary information from other variables.
@@
In Section 2, we develop our FABLE methodology with the CC-FABLE modification and provide an approach to choose key hyperparameter values. In Section 3, we provide theoretical results on pseudo-posterior contraction rates and uncertainty quantification, leveraging on a blessing of dimensionality phenomenon. In Section 4, we validate our approach with comparison to other state-of-the-art approaches from the viewpoint of estimation error, frequentist coverage, and computational runtime in different numerical experiments. In Section 5, we present a relevant application of our approach to a large gene expression data set and highlight distinct advantages over competitors. In Section 6, we provide a discussion on future research directions.
@@ 2.1 Initial Approach
where we have omitted the intercept term, assuming the data have been centered prior to analysis. Here, $\Lambda\in\mathbb{R}^{p\times k}$ is an unknown matrix of factor loadings, $\eta_{i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$ is the $i$th latent factor, and $\epsilon_{i}$ is a zero-mean idiosyncratic error having diagonal covariance $\Sigma=\mbox{diag}(\sigma_{1}^{2},\ldots,\sigma_{p}^{2})$ for $i=1,\ldots,n$. It is typical to assume that the dimension of the latent factors $k$ is much smaller than the dimension of the data $p$. Integrating out the latent factors leads to
@@
The covariance matrix $\Psi=\Lambda\Lambda^{\top}+\Sigma$ of the $y_{i}$s in (2) is decomposed as the sum of a low-rank and a diagonal matrix. In this paper, our goal is to estimate $L=\Lambda\Lambda^{\top},\Sigma,$ and $\Psi=\Lambda\Lambda^{\top}+\Sigma$. We first illustrate our methodology assuming $k$ is known. We discuss an approach to estimate $k$ in Section 2.3.
@@
where $\epsilon^{(j)}$ is the $j$th column of the matrix $E=[\epsilon_{1},\ldots,\epsilon_{n}]^{\top}=[\epsilon^{(1)},\ldots,\epsilon^{(p)}],$ with $\epsilon^{(j)}\overset{ind}{\sim}N_{n}(0_{n},\sigma_{j}^{2}\mathbb{I}_{n})$ for $j=1,\ldots,p$. Writing in matrix form, we obtain
@@
Our strategy to obtain computationally efficient pseudo-posterior samples for $\Psi$ is outlined as follows. We first estimate $\mathbf{M}$ by a plug-in estimator $\widehat{\mathbf{M}}$ based on the singular value decomposition (SVD) of $\mathbf{Y}$ and then consider a surrogate model to (3) where $\mathbf{M}$ is replaced by $\widehat{\mathbf{M}}$. Conditional on the observed data, $\widehat{\mathbf{M}}$ is fixed. Thus, the surrogate model is equivalent to $p$ parallel regression problems. We then let the regression coefficient and the error variance for the $j$th regression have conjugate normal-inverse gamma prior distributions for $j=1,\ldots,p$ to obtain their pseudo-posterior samples across the $p$ different regressions in parallel. Finally, we combine the pseudo-posterior samples of the regression coefficients and the error variances to obtain a pseudo-posterior sample of $\Psi$.
To motivate our choice for $\widehat{\mathbf{M}}$, we start with the SVD of $\mathbf{Y}$, given by
@@
where $U\in\mathbb{R}^{n\times k},D\in\mathbb{R}^{k\times k},V\in\mathbb{R}^{p\times k},U_{\perp}\in\mathbb{R}^{n\times(r-k)},D_{\perp}\in\mathbb{R}^{(r-k)\times(r-k)},$ and $V_{\perp}\in\mathbb{R}^{p\times(r-k)}$, with $r=p\wedge n.$ The columns of $U,U_{\perp},V,V_{\perp}$ consist of orthonormal vectors, with $U^{\top}U_{\perp}=V^{\top}V_{\perp}=\mathbf{O}_{k\times(r-k)}$, where $\mathbf{O}_{p_{1}\times p_{2}}$ denotes the $p_{1}\times p_{2}$ matrix with all entries equal to 00. Let $\mathbf{A}=\mathbf{Y}V/\sqrt{p}=UD/{\sqrt{p}}$ and $\widehat{C}\in\mathbb{R}^{k\times k}$ be an invertible matrix satisfying
@@
For the choice of $\widehat{C}=D/\sqrt{np},$ it is straightforward from (7) that $\widehat{\mathbf{M}}=\sqrt{n}U.$ We will denote this particular choice of $\widehat{\mathbf{M}}=\sqrt{n}U$ as the canonical choice of $\widehat{\mathbf{M}}$, observing that other choices are also possible for different choices of $\widehat{C}$ satisfying (6). We provide a heuristic leading to (7) at the end of this Section.
@@
for $j=1,\ldots,p$. The model (8) could be interpreted as a version of (3) with $\widehat{\mathbf{M}}$ substituted for the original matrix of latent factors $\mathbf{M}$ and new parameters $\widetilde{\lambda}_{j}\in\mathbb{R}^{k}$, $\widetilde{\sigma}_{j}^{2}>0$ for $j=1,\ldots,p$. Next, we endow $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})_{j=1}^{p}$ with independent normal-inverse gamma (NIG) priors $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})\overset{iid}{\sim}\mbox{NIG}(0_{k},\tau^{2}\mathbb{I}_{k},\gamma_{0}/2,\gamma_{0}\delta_{0}^{2}/2)$. That is, we let
@@
The global shrinkage parameter $\tau^{2}$ allows us to a priori shrink the factor loadings towards zero, regularizing $\widetilde{\Lambda}=[\widetilde{\lambda}_{1},\ldots,\widetilde{\lambda}_{p}]^{\top}$ and favoring sparsity. We discuss a data-driven strategy to estimate $\tau^{2}$ in Section 2.3.
The surrogate model (8) and the conjugate prior specification (9) motivate the pseudo-posterior densities $\widetilde{\Pi}_{j}$ of $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})$ for each $j=1,\ldots,p$, given by
@@
$\begin{split}\widetilde{\Pi}_{j}(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})&=\mbox{NIG}\left(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2}\mid\mu_{j},\mathbf{K},\gamma_{n}/2,\gamma_{n}\delta_{j}^{2}/2\right)\\ &=N_{k}\left(\widetilde{\lambda}_{j}\mid\mu_{j},\widetilde{\sigma}_{j}^{2}\mathbf{K}\right)\mbox{IG}\left(\widetilde{\sigma}_{j}^{2}\,\middle|\,\dfrac{\gamma_{n}}{2},\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\right),\end{split}$ (10)
with the updated hyperparameters given by
@@
$\begin{split}\mu_{j}&=\left(\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+\dfrac{\mathbb{I}_{k}}{\tau^{2}}\right)^{-1}\widehat{\mathbf{M}}^{\top}y^{(j)},\\ \mathbf{K}&=\left(\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+\dfrac{\mathbb{I}_{k}}{\tau^{2}}\right)^{-1},\\ \gamma_{n}&=\gamma_{0}+n,\\ \gamma_{n}\delta_{j}^{2}&=\gamma_{0}\delta_{0}^{2}+\left(y^{(j)\top}y^{(j)}-\mu_{j}^{\top}\mathbf{K}^{-1}\mu_{j}\right).\end{split}$ (11)

### v3

@@ Abstract
Bayesian factor analysis is routinely used for dimensionality reduction in modeling of high-dimensional covariance matrices. Factor analytic decompositions express the covariance as a sum of a low rank and diagonal matrix. In practice, Gibbs sampling algorithms are typically used for posterior computation, alternating between updating the latent factors, loadings, and residual variances. In this article, we exploit a blessing of dimensionality to develop a provably accurate posterior approximation for the covariance matrix that bypasses the need for Gibbs or other variants of Markov chain Monte Carlo sampling. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach relies on a first-stage singular value decomposition (SVD) to estimate the latent factors, and then defines a jointly conjugate prior for the loadings and residual variances. The accuracy of the resulting posterior approximation for the covariance improves with increasing samples as well as increasing dimensionality. We show that FABLE has excellent performance in high-dimensional covariance matrix estimation, including producing well-calibrated credible intervals, both theoretically and through simulation experiments. We also demonstrate the strength of our approach in terms of accurate inference and computational efficiency by applying it to a gene expression dataset.
@@ 1 Introduction
Inference on covariance matrices in high-dimensional data is a key focus in many application areas, motivating a rich literature on associated statistical methods. One thread of this literature focuses on frequentist estimators of high-dimensional covariance matrices. Such approaches avoid direct modeling of the data and simply estimate the covariance under various assumptions on its inherent low-dimensional structure, including (but not limited to) banded covariance (Bickel and Levina, 2008b, ), low rank structure (Shikhaliev et al.,, 2019), low rank with sparsity (Richard et al.,, 2012), sparse covariance (Bien and Tibshirani,, 2011), and sparse inverse precision matrix estimation (Zhang and Zou,, 2014). Our interest is instead in model-based Bayesian approaches, which have advantages in terms of naturally accommodating complexities in the data and quantifying uncertainty, while having disadvantages in terms of computational efficiency, particularly in implementing Markov chain Monte Carlo (MCMC) algorithms to sample from the posterior distribution.
@@
Our particular focus is on factor models, which express the covariance matrix as a sum of low rank and diagonal matrices, and represent a popular probabilistic model-based alternative to principal components analysis (PCA). Factor analysis introduces a lower-dimensional latent factor $\eta_{i}\in\mathbb{R}^{k}$ corresponding to each observation $y_{i}\in\mathbb{R}^{p}$ with $k\ll p$, such that $y_{i}=\Lambda\eta_{i}+\epsilon_{i}$ for a factor loadings matrix $\Lambda\in\mathbb{R}^{p\times k}$ and independent idiosyncratic errors $\epsilon_{i}\sim N_{p}(0,\Sigma)$ with diagonal $\Sigma$. Following typical practice, we let the latent factors have independent Gaussian priors $\eta_{i}\sim N_{k}(0,\mathbb{I}_{k})$. Marginalizing out the latent factors leads to $y_{i}\sim N_{p}(0,\Psi)$, with the covariance matrix $\Psi=\Lambda\Lambda^{\top}+\Sigma$. This provides a convenient decomposition of the covariance, which dramatically reduces the number of free parameters required to model the $p\times p$ covariance matrix from $\mathcal{O}(p^{2})$ to $\mathcal{O}(pk)$, since $\Lambda\in\mathbb{R}^{p\times k}$ and $\Sigma\in\mathbb{R}^{p\times p}$ is diagonal.
@@
There is a vibrant recent literature improving upon and expanding the scope of Bayesian factor analysis (Schiavon et al.,, 2022; De Vito et al.,, 2021; Frühwirth-Schnatter,, 2023; Roy et al.,, 2021; Ma and Liu,, 2022; Bolfarine et al.,, 2022; Xie et al.,, 2022). Even with increasingly rich classes of priors and data types such as massive binary/count data, the canonical approach for posterior computation remains Gibbs samplers that iterate between updating the latent factors, factor loadings, residual variances, hyperparameters controlling the hierarchical prior, and other model parameters. This approach is simple to implement in broad model classes, but commonly faces problems with slow mixing, particularly as data dimensionality and complexity increase. Issues with slow mixing and bottlenecks with large datasets have motivated alternative posterior sampling algorithms that apply Hamiltonian Monte Carlo (HMC) after marginalizing out the latent factors, exploiting sufficient statistics and parallel computation to facilitate implementation for large sample sizes (Chandra et al.,, 2025; Man and Culpepper,, 2022). Nonetheless, all current MCMC sampling algorithms for large covariance matrices face substantial computational hurdles as the number of dimensions increases.
@@
The lack of sufficiently scalable MCMC algorithms has motivated a rich literature on developing more computationally efficient algorithms for Bayesian inference in factor models. Variational Bayes approaches (Hansen et al.,, 2025; Wang and Stephens,, 2021) attempt to approximate the posterior but often the accuracy of the approximation is poor, with a tendency to dramatically underestimate posterior uncertainty. A concise review comparing variational approximations and posterior sampling approaches for factor analysis is provided in Foo and Shim, (2021). Maximum a posteriori (MAP) estimation under sparsity priors (Srivastava et al.,, 2017) and expectation-maximization (EM) algorithms (Ročková and George,, 2016; Avalos-Pacheco et al.,, 2022; Zhao et al.,, 2016) provide computationally efficient point estimates of the covariance matrix. Sabnis et al., (2016) develop a distributed computing framework aggregating local estimates of the covariance across different cores. Srivastava et al., (2017), Ročková and George, (2016), and Sabnis et al., (2016) provide theoretical results regarding estimation accuracy and posterior concentration, without providing any guarantees on uncertainty quantification (UQ). A key open problem in the literature remains how to obtain an efficient posterior approximation for a high-dimensional covariance that has guarantees in terms of valid frequentist coverage of credible intervals.
@@
Fast algorithms for Bayesian factor analysis capable of scaling efficiently to high-dimensional data typically sacrifice the ability to provide an accurate characterization of uncertainty. The focus of this article is on proposing a simple approach for overcoming this limitation, providing computationally efficient inference on a high-dimensional covariance matrix. Assuming the rank $k$ remains fixed, we have more and more variables loading on the same small number $k$ of latent factors with increasing dimensions. As a consequence, we obtain a blessing of dimensionality phenomenon allowing us to first pre-estimate the latent factors and then leverage this point estimate to obtain an approximate posterior distribution of the factor loadings $\Lambda$ and error variances $\Sigma$ under conjugate priors on these parameters. This in turn induces a posterior approximation for the covariance matrix $\Psi=\Lambda\Lambda^{\top}+\Sigma$. The error in estimating the covariance based on this posterior approximation decreases as both the number of samples and number of dimensions increases. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach completely bypasses MCMC. Furthermore, we develop a data-adaptive coverage-correction methodology which ensures that the credible intervals for the entries of the covariance have guarantees regarding valid frequentist coverage.
@@
Fan et al., (2024) also pre-estimate latent factors, but in the context of regression with high-dimensional correlated covariates. In their approach, a first-stage principal components analysis is carried out for the high-dimensional covariates. The results are then used to obtain a modified covariate vector consisting of principal component scores and de-correlated versions of the original covariates. Using this new covariate vector in the second stage high-dimensional regression avoids multicollinearity. This focus, and the corresponding methodology, are substantially different from ours. A blessing of dimensionality phenomenon is also explored in Li et al., (2018), where estimation of the covariance between a subset of the variables of interest is improved by incorporating auxiliary information from other variables. Their methodology is also considerably different from ours, although we demonstrate similar gains with FABLE in a real-world gene expression data application when estimating covariance submatrices of relevant subsets of genes.
@@
In Section 2, we develop our FABLE methodology and provide an approach to choose key hyperparameter values. In Section 3, we provide theoretical results on posterior contraction rates and uncertainty quantification, leveraging a blessing of dimensionality phenomenon. In Section 4, we validate our approach through comparison with other state-of-the-art approaches from the viewpoint of estimation error, frequentist coverage, and computational runtime in various numerical experiments. In Section 5, we present a relevant application of our approach to a large gene expression dataset and highlight distinct advantages over competitors. In Section 6, we provide a discussion on future research directions.
@@ 2.1 Overview
where we have omitted the intercept term, assuming the data have been centered prior to analysis. Here, $\Lambda\in\mathbb{R}^{p\times k}$ is an unknown matrix of factor loadings, $\eta_{i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$ is the $i$th latent factor, and $\epsilon_{i}$ is a zero-mean idiosyncratic error having diagonal covariance $\Sigma=\mbox{diag}(\sigma_{1}^{2},\ldots,\sigma_{p}^{2})$ for $i=1,\ldots,n$. Integrating out the latent factors leads to
@@
In this paper, our goal is to estimate $L=\Lambda\Lambda^{\top},\Sigma,$ and $\Psi=\Lambda\Lambda^{\top}+\Sigma$. It is typical to assume that $k\ll p$. This facilitates dimension reduction of the high-dimensional $y_{i}\in\mathbb{R}^{p}$ to the lower-dimensional $\eta_{i}\in\mathbb{R}^{k}$, along with ensuring that the high-dimensional covariance matrix $\Psi\in\mathbb{R}^{p\times p}$ of the $y_{i}$s may be interpreted as the sum of low-rank and diagonal matrices. We first illustrate our methodology assuming $k$ is known and later discuss an approach to estimate $k$ in Section 2.4.1.
@@
where $\epsilon^{(j)}$ is the $j$th column of the matrix $E=[\epsilon_{1},\ldots,\epsilon_{n}]^{\top}=[\epsilon^{(1)},\ldots,\epsilon^{(p)}],$ with $\epsilon^{(j)}\overset{ind}{\sim}N_{n}(0,\sigma_{j}^{2}\mathbb{I}_{n})$ for $j=1,\ldots,p$. Writing in matrix form, we obtain
@@
We now outline our strategy to obtain an approximation for the posterior of $\Psi$ from which computationally efficient samples can be drawn. First, we estimate $\mathbf{M}$ by a plug-in estimator $\widehat{\mathbf{M}}$ based on the singular value decomposition (SVD) of $\mathbf{Y}$ and then consider a surrogate model to (3) where $\mathbf{M}$ is replaced by $\widehat{\mathbf{M}}$. Conditional on the observed data, $\widehat{\mathbf{M}}$ is fixed. Thus, the surrogate model is equivalent to $p$ parallel regression problems. Next, in the surrogate model, we endow the regression coefficient and the error variance for the $j$th regression with conjugate normal-inverse gamma prior distributions for $j=1,\ldots,p$. We then sample the regression coefficients and error variances jointly from the posterior approximations, across the $p$ different regressions in parallel. Finally, we combine the obtained samples of the regression coefficients and the error variances to obtain a sample of $\Psi$ drawn from the posterior approximation.
To illustrate our choice for $\widehat{\mathbf{M}}$, we start with the SVD of $\mathbf{Y}$, given by
@@
where $U\in\mathbb{R}^{n\times k},D\in\mathbb{R}^{k\times k},V\in\mathbb{R}^{p\times k},U_{\perp}\in\mathbb{R}^{n\times(r-k)},D_{\perp}\in\mathbb{R}^{(r-k)\times(r-k)},$ and $V_{\perp}\in\mathbb{R}^{p\times(r-k)}$, with $r=n\wedge p.$ The columns of $U,U_{\perp},V,V_{\perp}$ consist of orthonormal vectors, with $U^{\top}U_{\perp}=V^{\top}V_{\perp}=\mathbf{O}_{k\times(r-k)}$, where $\mathbf{O}_{p_{1}\times p_{2}}$ denotes the $p_{1}\times p_{2}$ matrix with all entries equal to 0. Let $\mathbf{A}=\mathbf{Y}V/\sqrt{p}=UD/{\sqrt{p}}$ and $\widehat{C}\in\mathbb{R}^{k\times k}$ be an invertible matrix satisfying
@@
For the choice of $\widehat{C}=D/\sqrt{np},$ it is straightforward from (7) that $\widehat{\mathbf{M}}=\sqrt{n}U.$ We will denote this particular choice of $\widehat{\mathbf{M}}=\sqrt{n}U$ as the canonical choice of $\widehat{\mathbf{M}}$. However, other choices of $\widehat{\mathbf{M}}$ are also possible for different choices of $\widehat{C}$ satisfying (6), such as $\widehat{C}=DQ/\sqrt{np}$ for any $k\times k$ rotation matrix $Q$ satisfying $QQ^{\top}=\mathbb{I}_{k}$. If either $n$ or $p$ is very large, approximate SVD approaches (Halko et al.,, 2011) could be implemented instead of (5) to improve computational efficiency. We provide a heuristic leading to (7) in Section 2.2.
@@
for $j=1,\ldots,p$. The model (8) could be interpreted as a version of (3) with $\widehat{\mathbf{M}}$ substituted for the original matrix of latent factors $\mathbf{M}$ and surrogate model parameters $\widetilde{\lambda}_{j}\in\mathbb{R}^{k}$, $\widetilde{\sigma}_{j}^{2}>0$ for $j=1,\ldots,p$, endowed with independent normal-inverse gamma (NIG) priors $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})\overset{iid}{\sim}\mbox{NIG}(0_{k},\tau^{2}\mathbb{I}_{k},\gamma_{0}/2,\gamma_{0}\delta_{0}^{2}/2)$. That is, we let
@@
The global shrinkage parameter $\tau^{2}$ allows us to a priori shrink the factor loadings towards zero, regularizing $\widetilde{\Lambda}=[\widetilde{\lambda}_{1},\ldots,\widetilde{\lambda}_{p}]^{\top}$. We discuss a data-driven strategy to estimate $\tau^{2}$ in Section 2.4.
The surrogate model (8) and the conjugate prior specification (9) motivate the sampling of
@@
$\widetilde{\lambda}_{j}\mid\widetilde{\sigma}_{j}^{2}\sim N_{k}(\mu_{j},\rho^{2}\widetilde{\sigma}_{j}^{2}\mathbf{K}),\quad\widetilde{\sigma}_{j}^{2}\sim\mbox{IG}(\gamma_{n}/2,\gamma_{n}\delta_{j}^{2}/2),\quad\text{for $j=1,\ldots,p$, where}$ (10)
@@
$\hskip-5.0pt\begin{array}[]{cc}\displaystyle\begin{aligned} \mathbf{K}&=\left(\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+\frac{\mathbb{I}_{k}}{\tau^{2}}\right)^{-1}&=\frac{\mathbb{I}_{k}}{n+\tau^{-2}},\\[10.00002pt] \mu_{j}&=\left(\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+\frac{\mathbb{I}_{k}}{\tau^{2}}\right)^{-1}\widehat{\mathbf{M}}^{\top}y^{(j)}&=\frac{\widehat{\mathbf{M}}^{\top}y^{(j)}}{n+\tau^{-2}},\end{aligned}&\displaystyle\begin{aligned} \gamma_{n}&=\gamma_{0}+n,\\[10.00002pt] \gamma_{n}\delta_{j}^{2}&=\gamma_{0}\delta_{0}^{2}+\left(y^{(j)\top}y^{(j)}-\mu_{j}^{\top}\mathbf{K}^{-1}\mu_{j}\right),\end{aligned}\end{array}$ (12)
since from Proposition 2.1 in Section 2.2, we have $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}=n\,\mathbb{I}_{k}$ for any $\widehat{\mathbf{M}}$ satisfying (7). We denote this posterior approximation by $\widetilde{\Pi}_{j}$, namely
$\begin{split}\widetilde{\Pi}_{j}(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})&=\mbox{NIG}\left(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2}\mid\mu_{j},\rho^{2}\mathbf{K},\gamma_{n}/2,\gamma_{n}\delta_{j}^{2}/2\right)=N_{k}\left(\widetilde{\lambda}_{j}\mid\mu_{j},\rho^{2}\widetilde{\sigma}_{j}^{2}\mathbf{K}\right)\mbox{IG}\left(\widetilde{\sigma}_{j}^{2}\,\middle|\,\dfrac{\gamma_{n}}{2},\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\right),\end{split}$ (13)
The posteriors $\widetilde{\Pi}_{j}$ are simply obtained by treating $\widehat{\mathbf{M}}$ as a fixed design matrix for the $j$th surrogate regression and then applying Bayes’ rule. The additional coefficient $\rho^{2}$ when drawing the $\widetilde{\lambda}_{j}$s conditional on the sampled $\widetilde{\sigma}_{j}^{2}$ ensures the correct average asymptotic coverage of entrywise Bayesian credible intervals. That is, the average asymptotic frequentist coverage of the $100(1-\alpha)\%$ credible intervals corresponding to the entries of the covariance matrix $\widetilde{\Psi}$ is also $1-\alpha$. We describe the strategy to appropriately choose $\rho^{2}$ in Section 2.3.